# LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning

### v2

@@ Abstract
Keywords: Hierarchical Reinforcement Learning (HRL), Large Language Models (LLM), Robot Navigation, Robot Manipulation
@@ 1 Introduction
Figure 1: LGR2 overview (left): The higher-level policy predicts subgoals $g_{t}$ for the lower-level policy, which executes primitive actions $a_{t}$ on the environment. The lower-level replay buffer is populated by environment interactions, and RL is subsequently used to optimize the lower-level policy. L2R is used to translate human instructions to reward function parameters, and is subsequently used to relabel higher-level replay buffer transitions. Finally, RL is used to optimize the higher-level policy.
@@
While deep reinforcement learning (RL) has achieved notable advancements in executing complex robotic tasks [1, 2, 3, 4], its success is often hindered by challenges such as ineffective exploration and long-term credit assignment, especially in sparse-reward scenarios [5]. Hierarchical reinforcement learning (HRL) [6, 7, 8, 9, 10] presents an elegant framework that offers the benefits of temporal abstraction and enhanced exploration [11] to address these issues. In goal-conditioned RL setting [7, 8], the lower-level policy executes primitive actions to achieve the subgoals provided by the higher-level policy. However, the performance in off-policy HRL is often impeded by issues like non-stationarity, and therefore, developing more sophisticated techniques for solving HRL are essential.
Large Language Models (LLMs) trained on internet data [12] have shown remarkable capabilities to solve complex tasks ranging from logical reasoning [13, 14] and code generation [15]. Considerable advancement have also been made in utilizing LLMs to guide robotic behaviours [16, 17, 18]. Using natural language instructions to directly solve complex robotic tasks is a crucial step towards building interactive robots. Recently, language instructions have been employed to generate parameters for reward function in L2R [19], which is subsequently optimized using MPC controllers [20]. This raises the following question: can we utilize reward generation capabilities of LLMs to build an HRL approach, that leverages natural language instructions to solve complex robotic control tasks, while simultaneously addressing the non-stationarity issue in HRL?
In this work, we provide an affirmative answer to the above question by proposing LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning. In LGR2, we first translate language instructions to reward function parameters using L2R approach, and then subsequently leverage the reward function parameters to relabel higher-level replay buffer transitions. The key insight behind the HRL component of LGR2 is that the higher-level policy is decoupled from the non-stationary lower-level primitive, which effectively alleviates the non-stationarity problem encountered in off-policy HRL. Additionally, since the relabeled rewards might be too sparse for the lower primitive to show any significant progress, we employ hindsight experience replay [21] to densify the higher-level rewards. To the best of our knowledge, ours is the first approach that effectively combines language-guided reward generation and HRL to learn effective policies for solving complex robotic control tasks. In order to assess the efficacy of the proposed method, we perform experiments across a range of complex robotic tasks. Our experiments demonstrate that DIPPER achieves remarkable performance and consistently outperforms existing baselines. We provide LGR2 overview in Figure 1.
We summarize the main contributions of LGR2 as follows:
(i) we propose a novel HRL framework for solving robotic control tasks from language instructions (Section 4),
(ii) we show that LGR2 is able to mitigate non-stationarity in off-policy HRL (Section 5),
(iii) we utilize hindsight relabeling to enhance sample efficiency and address the issue of sparsity in scenarios with sparse rewards, (Section 4.2), and
@@ 2 Related Work
Hierarchical Reinforcement Learning The development of effective hierarchies of policies has attracted considerable research interest in reinforcement learning (RL) community[22, 23, 24, 25]. The options framework[23, 26, 27, 28, 29, 30] focuses on learning temporally extended macro actions and termination functions to tackle long-horizon tasks. However, without proper regularization, these methods can often result in degenerate solutions. In goal-conditioned learning, previous approaches have narrowed the search space by solving for specific goals in a greedy manner [31, 32], and this strategy has been adapted to hierarchical RL (HRL) as well [33, 34, 35]. Methods such as HIRO [36] and HAC [37] address the non-stationarity issue in hierarchical learning by relabeling subgoals within the replay buffer. In contrast, LGR2 addresses non-stationarity in HRL by utilizing LLMs to define reward parameters, which are optimized using HRL to solve complex robotic control tasks.
@@
Other works have attempted to pre-train policies on related tasks to develop behaviour skill priors [38, 39], followed by fine-tuning using RL. However, these approaches may struggle to generalize when there is a distributional shift between the initial and target task distributions, and they may also fail to learn effective policies due to sub-optimal priors. Some methods rely on hand-designed action primitives [40, 41] to determine how to solve sub-tasks and subsequently predict what to do by selecting the appropriate primitives. The design of these action primitives can be labour-intensive, limiting the applicability of such approaches to complex real-world tasks. LGR2 avoids these limitations by learning hierarchical policies in parallel, thus eliminating the need for hand-designed action primitives and enhancing the ability to handle more complex tasks.
@@
Language to Actions The earliest work of this paradigm studied mapping templated language to controllers with temporal logic [42] or learning a parser to motion primitives [43]. Recent studies have focussed on utilizing end-to-end models that produce actions based on natural language descriptions.  Ku et al. [44] have used natural language descriptions for instructions for the navigation methods of robots. However, they suffer from the assumption of low-dimensional actions navigating from one graph node to another [44, 45]. Utilization of latent embeddings of language commands as multiple input text is a commonly used approach to extend the end-to-end approaches for navigation manipulation. These embeddings are either trained with behavioural cloning [46, 47, 48], offline reinforcement learning [49], goal-conditioned reinforcement learning [50], or in a shared autonomy paradigm [51] to translate the natural language descriptions into robot navigation methods. End-to-end policies can be very effective, but they require a large amount of training data. [19] proposed a less data-centric method in which an optimal controller generates low-level actions. In contrast, we employ an HRL based approach that leverages temporal abstraction to induce training efficiency.
@@
Language to Code Large Language Models (LLMs) like Llama [52], GPT-4 [53], and Gemini [54] have been extensively used for code generation both in and outside the robotics context [15, 55, 56]. These models are highly efficient in solving coding competition questions [57], drawing figures [58], generating policies to solve 2D tasks [59], and complex instruction tasks [16]. In this work, we use LLMs to generate language-guided higher-level rewards.
@@
Language to Rewards Several works have explored the idea of translating natural language instructions to rewards [60, 61, 62, 63, 64, 65, 66]. Prior works have gone into training domain-specific reward models to map language instructions to reward values [60, 61, 62] or constraints [60]. These models can perform efficiently for challenging robotic tasks such as object pushing [60] and drawer opening [62] but are hindered by the requirement of a large amount of language-labelled data. Recent research has focused on employing LLMs as a reward function to infer user intents in negotiation games or collaborative human-AI interaction games [65, 64]. All these works impeded utilizing the LLMs to assign reward values during RL training. Nevertheless, in training RL rules, these works get reward values of rollouts, necessitating a high volume of inquiries to LLMs. AutoRL [67] proposed an automated parameterization of reward functions, but they did not provide a language interface. We leverage reward generated from of language instructions to mitigate non-stationarity in HRL, and thus propose an effective framework for translating language instructions to robot actions.
@@ 3 Problem Formulation
In this paper, we explore the Markov decision process (MDP) framework denoted as $(\mathnormal{S},\mathnormal{A},p,r,\gamma)$. Here, $\mathnormal{S}$ represents the state space, and $\mathnormal{A}$ denotes the action space. The transition probability function, $p:\mathnormal{S}\times\mathnormal{A}\rightarrow\Delta(\mathnormal{S})$, maps state-action pairs to probability distributions over the state space. The reward function, $r:\mathnormal{S}\times\mathnormal{A}\rightarrow\mathnormal{R}$, assigns rewards based on state-action pairs, and $\gamma\in(0,1)$ is the discount factor. At each timestep $t$, the agent occupies state $s_{t}$ and selects an action $a_{t}$ based on a policy $\pi:\mathnormal{S}\rightarrow\Delta(\mathnormal{A})$, which maps states to probability distributions over the action space. The action $a_{t}$ is sampled from this distribution, $a_{t}\sim\pi(\cdot|s_{t})$. Following the action, the agent receives a reward $r_{t}=r(s_{t},a_{t})$, and the system transitions to a new state $s_{t+1}$, sampled according to the transition function, $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. In the traditional RL setup, we optimize the following objective: $\pi^{*}:=\arg\max_{\pi}J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right].$ In the following discussion, we will focus on the standard goal-conditioned setting as described in [21], where the agent’s policy is conditioned on both the current state and a specified goal. Specifically, at each timestep $t$, the policy $\pi$ determines actions $a_{t}$ based on both the current state $s_{t}$ and the desired goal $g_{t}$, such that $a_{t}\sim\pi(\cdot|s_{t},g_{t})$.
3.1 Hierarchical Reinforcement Learning
@@
In our goal-conditioned hierarchical framework, the higher-level policy sets subgoals for the lower-level policy to which in turn executes primitive actions on the environment. Specifically, the higher-level policy, denoted as $\pi^{H}:\mathnormal{S}\rightarrow\Delta(\mathnormal{G})$, selects subgoals $g_{t}$ from the set of possible goals $\mathnormal{G}\subset\mathnormal{S}$. At each timestep $t$, the higher-level policy predicts a subgoal $g_{t}\sim\pi^{H}(\cdot|s_{t})$ every $k$ timesteps, maintaining the subgoal as $g_{t}=g_{k\cdot\lceil t/k\rceil}$ in the intervals between these updates. Consequently, the higher-level policy updates the subgoals at regular intervals of $k$ timesteps, while the lower-level policy focuses on executing primitive actions to accomplish these subgoals in the meantime. Moreover, at each timestep $t$, the lower-level policy $\pi^{L}:\mathnormal{S}\times\mathnormal{G}\rightarrow\Delta(\mathnormal{A})$ selects primitive actions $a_{t}$ based on the current state $s_{t}$ and the subgoal $g_{t}$ provided by the higher-level policy. The action $a_{t}$ is sampled as $a_{t}\sim\pi^{L}(\cdot|s_{t},g_{t})$, leading to a state transition $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. The lower-level policy receives a reward $r^{L}_{t}=r^{L}(s_{t},g_{t},a_{t})$, defined as $-\mathbf{1}_{\{\|s_{t}-g_{t}\|_{2}>\varepsilon\}}$, where $\mathbf{1}_{G}$ is the indicator function for set $G$.
@@
In our goal-conditioned HRL framework, we represent the $k$-length trajectories of lower-level primitive behavior as sequences denoted by $\Delta$, defined as $\Delta=((s_{t},a_{t}),(s_{t+1},a_{t+1}),\ldots,(s_{t+k-1},a_{t+k-1}))$. Similarly, the $n$-length trajectories of higher-level behavior are represented as sequences denoted by $\sigma$, which consist of states and subgoal predictions, expressed as $\sigma=((s_{t},g_{t}),(s_{t+k},g_{t+k}),\ldots,(s_{t+(n-1)k},g_{t+(n-1)k}))$. In the vanilla HRL setup, where both levels are simultaneously trained using RL, the higher-level policy receives a reward $r^{H}_{t}=r^{H}(s_{t},g^{*},g_{t})$, where $g^{*}\in\mathnormal{G}$ is the ultimate goal and $r^{H}:\mathnormal{S}\times\mathnormal{G}\times\mathnormal{G}\rightarrow\mathbb{R}$ is the higher-level reward function. The lower-level policy stores samples in its replay buffer in the form $(s_{t},g_{t},a_{t},r^{L}_{t},s_{t+1})$. Meanwhile, after every $k$ timesteps, the higher-level policy stores samples of the form $(s_{t},g^{*},g_{t},r^{H}_{t}=\sum_{i=t}^{t+k-1}r^{L}_{i},s_{t+k})$ in its buffer, where $r^{H}_{t}$ is the sum of environment rewards encountered by the lower policy for $k$ timesteps.
3.1.1 Limitations of standard HRL approaches
@@
Although HRL offers notable benefits over traditional non-hierarchical RL, including enhanced sample efficiency through temporal abstraction and improved exploration [68, 11], it still suffers from serious limitations. As highlighted in [68] and [69], off-policy HRL suffers from non-stationarity due to the non-stationary lower-level policy behavior. Specifically, the transitions in the higher-level replay buffer collected using previous versions of the lower-level policy, become outdated as the lower-level policy changes. Consequently, despite the theoretical advantages of HRL, it often performs poorly in practice [68, 69, 70]. In this work, we focus on developing a novel HRL approach that leverages the advancements in LLMs to mitigate the recurring issue of non-stationarity in HRL.
3.2 Language to Rewards
@@
In prior work [60, 71, 65], natural language instructions have been translated to rewards using domain specific reward models. In this work, we consider the work L2R [19], which uses the language instructions to generate reward parameters. These reward parameters act as an interface between high-level language instructions and lower-level primitive actions to accomplish a variety of robotic tasks. In L2R, a reward translator module translates the natural language instructions into reward parameters, which are further optimized using Mujoco MPC controller [20]. In this work, instead of the MPC controller, we use RL to optimize higher and lower level policies. The reward translator consists of two pre-defined templates: the motion descriptor and the reward coder, which are shown to successfully generate reward parameters for complex tasks. However, it faces two limitations:
@@
L1: In complex long horizon tasks, learning a single-level policy may prove overly challenging.
@@
L2: The generated reward function might be too sparse for any significant lower primitive progress.
To resolve these limitations, our novel LGR2 approach harnesses the benefits of temporal abstraction provided by HRL to deal with L1, and employs hindsight relabeling to deal with L2.

### v3

@@ 1 Introduction
While deep reinforcement learning (RL) has achieved notable advancements in executing complex robotic tasks (Rajeswaran et al., 2017; Kalashnikov et al., 2018; Gu et al., 2016; Levine et al., 2015), its success is often impeded by challenges like ineffective exploration and long-term credit assignment, especially in sparse-reward scenarios (Nair et al., 2017). To address such issues, goal-conditioned hierarchical reinforcement learning (HRL) (Sutton et al., 1999b; Dayan & Hinton, 1993; Vezhnevets et al., 2017; Klissarov et al., 2017b; Harb et al., 2018) offers the benefits of temporal abstraction and enhanced exploration (Nachum et al., 2019), where the lower-level policy executes primitive actions to achieve the subgoals provided by the higher-level policy. However, off-policy HRL suffers from non-stationarity due to changing lower level policy.
Large Language Models (LLMs) trained on internet data (Brown et al., 2020) have shown remarkable capabilities to solve complex tasks like logical reasoning (Wei et al., 2022; Kojima et al., 2022) and code generation (Chen et al., 2021). Considerable advancements have also been made in utilizing LLMs to solve effective robotic tasks (Liang et al., 2023; Huang et al., 2022; Ahn et al., 2022). Using natural language instructions to solve complex robotic tasks directly is a crucial step toward developing truly interactive robots. Recent studies have leveraged language instructions to parameterize reward functions in L2R frameworks (Yu et al., 2023) and subsequently optimizing these rewards with MPC controllers (Howell et al., 2022). Although these methods underscore the promise of language-guided control, they fall short when confronted with complex, long-term planning tasks, due to issues like long-term credit assignment and insufficient exploration, especially in sparse reward scenarios. Furthermore, while LLMs have internal knowledge of robot motions, they generally lack the specialized training data needed to generate direct, low-level control commands. The concept of grounding language to reward has also been explored in previous efforts to capture user preferences and task knowledge (Goyal et al., 2019; Lin et al., 2022b; Kwon et al., 2023). Despite promising outcomes, these typically rely on additional training data to learn the language-to-reward mapping. These observations naturally leads us to ask: can the reward generation capabilities of large language models be harnessed to develop an HRL framework that not only leverages natural language instructions for complex robotic control but also effectively addresses the non-stationarity inherent in HRL?
Figure 1: LGR2 overview (left): The higher-level policy predicts subgoals $g_{t}$ for the lower-level policy, which executes primitive actions $a_{t}$ on the environment. The lower-level replay buffer is populated by environment interactions, and reinforcement learning (RL) is subsequently used to optimize the lower-level policy. L2R is used to translate human instructions to reward function parameters, and is subsequently used to relabel higher-level replay buffer transitions. Finally, RL is used to optimize the higher-level policy.
@@
In this work, we provide an affirmative answer to the above question by proposing LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning. We propose a data-efficient interactive framework where users engage in dialogue with the LLM to guide the generation of rewards—thus steering robot behaviors without the need for extensive, task-specific datasets. Concretely, we first translate language instructions to reward function parameters, which are subsequently leveraged to relabel the higher-level replay buffer transitions. Since the higher-level policy is decoupled from the non-stationary lower-level primitive, this effectively mitigates the non-stationarity issue prevalent in off-policy HRL. Additionally, since the relabeled rewards might be too sparse for the lower primitive to show any significant progress, we employ hindsight experience replay (Andrychowicz et al., 2017) to densify the higher-level rewards. Thus, our approach effectively combines language-guided reward generation and HRL to solve complex robotic control tasks. To assess the efficacy of the proposed method, we perform experiments across a range of complex sparse robotic control tasks. Our experiments demonstrate that LGR2 achieves impressive performance and consistently outperforms prior baselines in various simulations and real-world tasks. We also empirically demonstrate that LGR2 mitigates non-stationarity in HRL, thereby stabilizing HRL training. We provide LGR2 overview in Figure 1.
We summarize our contributions as follows:
(i) We propose a novel HRL framework for solving robotic control tasks from language instructions (Section 4).
(ii) We show that LGR2 is able to mitigate non-stationarity in off-policy HRL (Section 5).
(iii) We utilize hindsight relabeling to enhance sample efficiency in sparse reward environments, (Section 4.2).
@@ 2 Related Work
Hierarchical Reinforcement Learning The development of effective hierarchies of policies has attracted considerable research interest in reinforcement learning (RL) community (Barto & Mahadevan, 2003; Sutton et al., 1999a; Parr & Russell, 1998; Dietterich, 1999). The options framework (Sutton et al., 1999a; Bacon et al., 2016; Harutyunyan et al., 2017; Harb et al., 2017; Harutyunyan et al., 2019; Klissarov et al., 2017a) focuses on learning temporally extended macro actions and termination functions to tackle long-horizon tasks. However, without proper regularization, these methods can often result in degenerate solutions. In goal-conditioned learning, previous approaches have narrowed the search space by solving for specific goals in a greedy manner (Kaelbling, 1993; Foster & Dayan, 2002), and this strategy has been adapted to hierarchical RL (HRL) as well (Wulfmeier et al., 2019; 2020; Ding et al., 2019). Further, off-policy HRL suffers from non-stationarity (Levy et al., 2018; Nachum et al., 2018). Prior works deal with these issues by either simulating optimal lower primitive behavior (Levy et al., 2018), relabeling replay buffer data (Nachum et al., 2018), assuming access to privileged information like expert demonstrations (Singh & Namboodiri, 2023b; a), or preferences (Singh et al., 2024b; Singh & Namboodiri, 2023b; Singh et al., 2024a). In contrast, LGR2 addresses non-stationarity in HRL by utilizing LLMs to define reward parameters, which are subsequently leveraged to solve complex robotic control tasks.
@@
Behavior Priors Other works have attempted to pre-train policies on related tasks to develop behaviour skill priors (Pertsch et al., 2020; Singh et al., 2020), followed by fine-tuning using RL. However, these approaches may struggle to generalize when there is a distributional shift between the initial and target task distributions, and may also fail to learn effective policies due to sub-optimal priors. Some methods rely on hand-designed action primitives (Dalal et al., 2021; Nasiriany et al., 2021) to determine how to solve sub-tasks and subsequently predict what to do by selecting the appropriate primitives, which can be labour-intensive, limiting the applicability of such approaches to complex real-world tasks. LGR2 avoids these limitations by learning hierarchical policies in parallel, thus eliminating the need for hand-designed action primitives and enhancing the ability to handle more complex tasks.
@@
Language to Actions The earliest work of this paradigm studied mapping templated language to controllers with temporal logic (Kress-Gazit et al., 2008) or learning a parser to motion primitives (Matuszek et al., 2012). Recent studies focus on utilizing end-to-end models that produce actions based on natural language descriptions.  (Ku et al., 2020) uses natural language descriptions for instructions for navigation methods of robots. However, they suffer from the assumption of low-dimensional actions navigating from one graph node to another (Ku et al., 2020; Kamath et al., 2023).
Utilizing latent embeddings of language commands as multiple input text is a commonly used approach to extend the end-to-end approaches for navigation manipulation. These embeddings are either trained with behavioural cloning (Mees et al., 2023; Jang et al., 2022; Lynch et al., 2022), offline RL (Ebert et al., 2021), goal-conditioned RL (Fu et al., 2019), or in a shared autonomy paradigm (Karamcheti et al., 2021) to translate the natural language descriptions into robot navigation methods. End-to-end policies require a large amount of training data. (Yu et al., 2023) propose a less data-centric approach where an optimal controller generates low-level actions. In contrast, we employ an HRL based approach that leverages temporal abstraction to induce training efficiency.
@@
Language to Code Large Language Models (LLMs) like Llama (Touvron et al., 2023), GPT-4 (OpenAI et al., 2024), and Gemini (Team et al., 2024) have been extensively used for code generation both in and outside the robotics context (Chen et al., 2021; Ellis et al., 2020; Austin et al., 2021). These models are highly efficient in solving coding competition questions (Li et al., 2022), drawing figures (Tian et al., 2020), generating policies to solve 2D tasks (Trivedi et al., 2022), and complex instruction tasks (Liang et al., 2023). This work uses LLMs to generate language-guided higher-level rewards for solving long horizon tasks.
@@
Language to Rewards Several works have explored the idea of translating natural language instructions to rewards (Sharma et al., 2022; Goyal et al., 2019; Nair et al., 2022; Bahdanau et al., 2018; Hu & Sadigh, 2023; Kwon et al., 2023; Lin et al., 2022a). Prior works have gone into training domain-specific reward models to map language instructions to reward values (Sharma et al., 2022; Goyal et al., 2019; Nair et al., 2022) or constraints (Sharma et al., 2022). These models can perform efficiently for challenging robotic tasks such as object pushing (Sharma et al., 2022) and drawer opening (Nair et al., 2022) but are hindered by the requirement of a large amount of language-labelled data. Recent research has focused on employing LLMs as a reward function to infer user intents in negotiation games or collaborative human-AI interaction games (Kwon et al., 2023; Hu & Sadigh, 2023). All these works impeded utilizing the LLMs to assign reward values during RL training. Nevertheless, in training RL rules, these works get reward values of rollouts, necessitating a high volume of inquiries to LLMs. AutoRL (Chiang et al., 2018) proposed an automated parameterization of reward functions, but they did not provide a language interface. In contrast, we leverage rewards generated from language instructions to propose an effective framework for directly translating language instructions to robot actions.
@@ 3 Problem Formulation
MDP Setup. The Markov decision process (MDP) framework is denoted as $(\mathnormal{S},\mathnormal{A},p,r,\gamma)$, where $\mathnormal{S}$ represents the state space, and $\mathnormal{A}$ denotes the action space. The transition probability function is denoted as $p:\mathnormal{S}\times\mathnormal{A}\rightarrow\Delta(\mathnormal{S})$. The reward function $r:\mathnormal{S}\times\mathnormal{A}\rightarrow\mathnormal{R}$, assigns rewards based on state-action pairs, and $\gamma\in(0,1)$ is the discount factor. At each timestep $t$, the agent occupies state $s_{t}$ and selects an action $a_{t}$ based on a policy $\pi:\mathnormal{S}\rightarrow\Delta(\mathnormal{A})$, which maps states to probability distributions over the action space. The action $a_{t}$ is sampled from this distribution, $a_{t}\sim\pi(\cdot|s_{t})$. Following the action, the agent receives a reward $r_{t}=r(s_{t},a_{t})$, and the system transitions to a new state $s_{t+1}$, sampled according to the transition function, $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. In the traditional RL setup, we optimize the following objective: $\pi^{*}:=\arg\max_{\pi}J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right].$ In the following discussion, we will focus on the standard goal-conditioned setting as described in (Andrychowicz et al., 2017), where the agent’s policy is conditioned on both the current state and a specified goal. Specifically, at each timestep $t$, the policy $\pi$ determines actions $a_{t}$ based on both the current state $s_{t}$ and the desired goal $g_{t}$, such that $a_{t}\sim\pi(\cdot|s_{t},g_{t})$.
@@
Hierarchical Reinforcement Learning. In our goal-conditioned hierarchical framework, the higher-level policy predicts subgoals for the lower-level policy which in turn executes primitive actions on the environment. Specifically, the higher-level policy, denoted as $\pi^{H}:\mathnormal{S}\rightarrow\Delta(\mathnormal{G})$, selects subgoals $g_{t}$ from the set of possible goals $\mathnormal{G}\subset\mathnormal{S}$. The higher-level policy predicts a subgoal $g_{t}\sim\pi^{H}(\cdot|s_{t})$ after every $k$ timesteps, and the lower-level policy $\pi^{L}:\mathnormal{S}\times\mathnormal{G}\rightarrow\Delta(\mathnormal{A})$ executes primitive actions $a_{t}$ to achieve the subgoals within the $k$ timestep window. The higher-level policy receives the environment reward $r^{H}_{t}=r^{H}(s_{t},g^{*},g_{t})$, where $g^{*}\in\mathnormal{G}$ is the ultimate goal and $r^{H}:\mathnormal{S}\times\mathnormal{G}\times\mathnormal{G}\rightarrow\mathbb{R}$ is the higher-level reward function which encourages progress toward the final goal objective. The lower-level policy receives the reward $r^{L}_{t}=-\mathbf{1}_{\{\|s_{t}-g_{t}\|_{2}>\varepsilon\}}$ which encourages progress towards the subgoal $g_{t}$, where $\mathbf{1}_{G}$ is the indicator function for set $G$.
@@
The $k$-length trajectories of lower-level primitive behavior are sequences denoted by $\Delta$, where $\Delta=((s_{t},a_{t}),(s_{t+1},a_{t+1}),\ldots,(s_{t+k-1},a_{t+k-1}))$. Similarly, the $n$-length trajectories of higher-level behaviour are sequences denoted by $\sigma$, which consist of states and subgoal predictions, expressed as $\sigma=((s_{t},g_{t}),(s_{t+k},g_{t+k}),\ldots,(s_{t+(n-1)k},g_{t+(n-1)k}))$. The lower-level policy stores samples in its replay buffer in the form $(s_{t},g_{t},a_{t},r^{L}_{t},s_{t+1})$. After every $k$ timesteps, the higher-level policy stores samples of the form $(s_{t},g^{*},g_{t},r^{H}_{t},s_{t+k})$ in its buffer.
@@
Limitations of standard HRL approaches. Although HRL offers notable benefits over traditional non-hierarchical RL, including enhanced sample efficiency through temporal abstraction and improved exploration (Nachum et al., 2018; 2019), it still suffers from serious limitations. As highlighted in Nachum et al. (2018) and Levy et al. (2018), off-policy HRL suffers from non-stationarity. Specifically, a high-level action may not yield the same low-level behavior in the future due to changing lower-level policy, thus invalidating that experience for training the high-level policy. Consequently, despite its theoretical advantages, HRL often performs poorly in practice (Nachum et al., 2018; Levy et al., 2018; Chane-Sane et al., 2021). In this work, we develop a novel approach that leverages the advancements in LLMs to mitigate the recurring issue of non-stationarity in HRL.
@@
Language to Rewards In prior work (Sharma et al., 2022; Lin et al., 2022b; Kwon et al., 2023), natural language instructions have been translated to reward functions using domain-specific reward models. For example, Yu et al. (2023) proposes a reward translator module consisting of two pre-defined templates: the motion descriptor and the reward coder to translate the natural language instructions to generate reward parameters. These parameters act as an interface between high-level language instructions and lower-level primitive actions.
Challenges of prior language to rewards (L2R) approaches.
  1. 1.
@@
    In complex long-horizon tasks, learning a single-level policy is challenging due to the issue of long-term credit assignment.
  2. 2.
@@
    The generated reward function might be too sparse for any significant lower primitive progress.
  3. 3.
    Prior works that leverage a pre-defined or manually-engineered library of control primitives to translate the natural language instructions into reward parameters require extensive expert knowledge or additional data collection.